# Simone Manoni*, Emanuele Parisi, Riccardo Tedeschi*, Davide Rossi* §, Andrea Acquaviva*, Andrea Bartolini*

### v1

@@ II-B CFI RISC-V ISA extension
Recently, RISC-V extended its ISA to support CFI through the introduction of Landing Pads and Shadow Stack extensions. The Zicfilp extension introduces a labeled landing pad instruction (lpad) encoded as auipc x0, label. The execution of an indirect jump forces the processor into a state where the only instruction allowed to retire next is a lpad instruction with a matching label, or a software-check exception is triggered. The expected label is pre-loaded in register x7 and must match the label encoded in the immediate field of the lpad instruction. The Zicfiss extension provides backward-edge protection via a shadow stack managed through dedicated instructions, including shadow stack push (sspush), pop and check (sspopchk), pointer read (ssrdp), and atomic swap (ssamoswap). Non-leaf functions save the link register into the shadow stack entry tracked by a dedicated shadow-stack pointer. Programs maintain both the regular and shadow stacks to protect return addresses from tampering. Unlike Intel’s CET, the return address is not transparently pushed and popped from the shadow stack in conjunction with function calls and returns.
@@ III-A Shadow Stack ISA extension design
Figure 1 illustrates the CVA6 pipeline, enhanced with the CFI extensions. The modules highlighted in blue have been introduced, or modified, to support the Zicfiss extension. The Control Status Registers (CSRs) have been extended to include the Shadow Stack pointer, and the environment configuration registers for machine, supervisor, and hypervisor modes have been integrated and expanded with a Shadow Stack enable field. This field is utilized in the frontend to compute the Shadow Stack-enabled state across various privilege levels and is then forwarded to the decode and execute stages. The decoder and compressed decoder are extended to handle the decoding of sspush, sspopchk, ssprd, and ssamoswap instructions, as well as their compressed counterparts. The sspush and sspopchk instructions have been mapped to store and load instructions, respectively, with special operation tags to ensure that Zicfiss operations are differentiated from standard memory operations in later pipeline stages. The same applies to the ssamoswap instructions, which are tagged with dedicated Shadow Stack swap W/D identifiers. We have implemented a dedicated Shadow Stack Unit (SSU) in the execution stage, responsible for performing shadow stack pop checks and early instruction filtering before allowing the instructions to access the LSU. If the SSU detects a ssamoswap operation executing in machine mode, or any Zicfiss instruction running with Supervisor or Virtual address translation and protection disabled, and at a privilege level below machine mode, a store access fault exception is raised and forwarded directly to the scoreboard. If neither of these two conditions is met, access to the LSU is allowed. Furthermore, when the SSU detects a sspopchk instruction, it buffers the link register content and the load transaction ID in dedicated registers and then waits for the load result. It also sets a flag to indicate that a sspopchk is in progress. When a sspopchk is in-flight, the SSU compares the load transaction ID of each load result coming from the LSU with the buffered transaction ID. If these IDs match, the SSU then checks the recorded link register value against the load result. If there is a mismatch between these values, a software check exception is raised. Additionally, we extended the MMU to enforce checks based on the type of page accessed. Zicfiss instructions can access only shadow stack pages, while non-Zicfiss can access only non-shadow stack pages. Any violation of this access policy results in a store access fault.

### v2

@@ II-B CFI RISC-V ISA extension
Recently, RISC-V extended its ISA to support CFI through the introduction of Landing Pads and Shadow Stack extensions. The Zicfilp extension introduces a labeled landing pad instruction (lpad) encoded as auipc x0, label. The execution of an indirect jump forces the processor into a state where the only instruction allowed to retire next is a lpad instruction with a matching label, or a software-check exception is triggered. The expected label is pre-loaded in register x7 and must match the label encoded in the immediate field of the lpad instruction. The Zicfiss extension provides backward-edge protection via a shadow stack managed through dedicated instructions, including shadow stack push (sspush), pop and check (sspchk), pointer read (ssrdp), and atomic swap (ssamoswap). Non-leaf functions save the link register into the shadow stack entry tracked by a dedicated shadow-stack pointer. Programs maintain both the regular and shadow stacks to protect return addresses from tampering. Unlike Intel’s CET, the return address is not transparently pushed and popped from the shadow stack in conjunction with function calls and returns.
@@ III-A Shadow Stack ISA extension design
Figure 1 illustrates the CVA6 pipeline, enhanced with the CFI extensions. The modules highlighted in blue have been introduced, or modified, to support the Zicfiss extension. The Control Status Registers (CSRs) have been extended to include the Shadow Stack pointer, and the environment configuration registers for machine, supervisor, and hypervisor modes have been integrated and expanded with a Shadow Stack enable field. This field is utilized in the frontend to compute the Shadow Stack-enabled state across various privilege levels and is then forwarded to the decode and execute stages. The decoder and compressed decoder are extended to handle the decoding of sspush, sspchk, ssprd, and ssamoswap instructions, as well as their compressed counterparts. The sspush and sspchk instructions have been mapped to store and load instructions, respectively, with special operation tags to ensure that Zicfiss operations are differentiated from standard memory operations in later pipeline stages. The same applies to the ssamoswap instructions, which are tagged with dedicated Shadow Stack swap W/D identifiers. We have implemented a dedicated Shadow Stack Unit (SSU) in the execution stage, responsible for performing shadow stack pop checks and early instruction filtering before allowing the instructions to access the load-store unit (LSU). If the SSU detects a ssamoswap operation executing in machine mode, or any Zicfiss instruction running with Supervisor or Virtual address translation and protection disabled, and at a privilege level below machine mode, a store access fault exception is raised and forwarded directly to the scoreboard. If neither of these two conditions is met, access to the LSU is allowed. Furthermore, when the SSU detects a sspchk instruction, it buffers the link register content and the load transaction ID in dedicated registers and then waits for the load result. It also sets a flag to indicate that a sspchk is in progress. When a sspchk is in-flight, the SSU compares the load transaction ID of each load result coming from the LSU with the buffered transaction ID. If these IDs match, the SSU then checks the recorded link register value against the load result. If there is a mismatch between these values, a software check exception is raised. Additionally, we extended the memory management unit (MMU) to enforce checks based on the type of page accessed. Zicfiss instructions can access only shadow stack pages, while non-Zicfiss can access only non-shadow stack pages. Any violation of this access policy results in a store access fault.